# When Debate Fails: Bias Reinforcement in Large Language Models

### v1

@@ 6.2 DReaMAD in Adversarial Strategic Decision-Making
Method NIM Fibonacci Kayles Chomp
Normal Misère Normal Misère Single 2 Rows Rectangular Square
GEMINI-1.5-flash
Standard Prompting 0.32 0.54 0.16 0.10 0.54 0.56 0.78 0.18
+ ReAct 0.10 0.68 0.16 0.76 0.50 0.30 0.42 0.12
+ Self-Refinement 0.14 0.66 0.18 0.36 0.50 0.46 0.46 0.16
+ Self-Consistency 0.04 0.28 0.28 0.86 - - - -
+ MAD 0.06 0.30 0.12 0.78 0.54 0.20 0.74 0.14
+ DReaMAD 0.38 0.84 0.16 0.94 0.58 0.62 0.60 0.22
GPT-4o-mini
Standard Prompting 0.38 0.54 0.22 0.28 0.46 0.48 0.46 0.38
+ ReAct 0.22 0.68 0.20 0.34 0.40 0.26 0.58 0.34
+ Self-Refinement 0.22 0.70 0.18 0.50 0.46 0.52 0.52 0.44
+ Self-Consistency 0.14 0.52 0.34 0.46 - - - -
+ MAD 0.28 0.62 0.22 0.82 0.42 0.28 0.52 0.56
+ DReaMAD 0.98 0.74 0.54 0.94 0.68 0.84 0.64 0.22
Due to technical issues, the self-consistency results for Kayles and Chomp are omitted and will be updated soon.
@@ 6.3 Does Generating Diverse Prompts Improve Performance?
Figure 7: Effect of Perspective Diversification by temperature. The Fibonacci dataset is evaluated 15 times with GPT-4o, averaged with a 95% confidence interval. Higher temperatures improve optimal action accuracy, showing that greater prompt diversity enhances reasoning performance.
We found that providing diverse prompts to multiple instances of the same model improves decision-making. To evaluate this, we generated identical prompts using the Strategic Prior Knowledge Elicitation module for two instances of the same model, denoted as (+)MAD, and distinct prompts for each instance, representing our proposed approach, as shown in Figure 6. We conducted experiments on four variants of MetaNIM Arena (NIM-Normal, NIM-Misère, Fibonacci-Normal, and Fibonacci-Misère) simulator (§A.8) total 50 episodes. Our results show that incorporating diverse prompts within the MAD framework significantly enhances performance, validating the effectiveness of our Perspective Diversification module.
Further, we examined the effect of varying temperature within the DReaMAD framework on the MetaNIM Arena Fibonacci dataset, adjusting it to 0.0, 0.2, 0.4, 0.6, 0.8, and 1.0 in the Strategic Prior Knowledge Elicitation and Perspective Diversification modules to assess its impact on prompt diversity. As shown in Figure 7, higher temperature (further diversity) correlates with increased optimal action accuracy, indicating that greater diversity in generated prompts contributes to improved reasoning performance.
@@ Appendix A MetaNIM Arena
Figure 8: The concept illustration of MetaNIM Arena. A detailed explanation of each game can be found in Appendix A.
@@
A.1 Sprague-Grundy Theorem
@@
A.2 Further Discussions on the Optimal Strategy
@@ A.6 Chomp
A.7 Our Dataset used in Experiments
@@
Table 5: Constructed dataset using MetaNIM Arena. We conduct an experiment on this dataset and report the results in Table 3.
@@
action space 1 - 3 dynamic (Max:30) x, y coordinate (differ to the scenario) a pin or adjacent two pins index
variants Normal Normal Square (2x2 $\sim$ 19x19) Normal
# of different samples 20 11 20 18
A.8 Our Simulator used in Experiments
@@
Features Kayles Chomp
Single 2 Rows Rectangular Square
Starting Point remaining items: 20 1 pile: 5, 2 pile: 6 2x8 5x5
Winning Condition taking last item taking last item avoiding poison avoiding poison
(top-left) (top-right)
@@
Action Space pile index pile index (row, column) x, y coordinate x, y coordinate
@@
Table 7: MetaNIM Arena Simulator: Kayles and Chomp
@@ D.1.2 Implications for Robust Prompting
Appendix E Ablation study setup

### v2

@@ 6.3 Does Generating Diverse Prompts Improve Performance?
We assess whether prompt diversity improves decision quality within the MAD framework. To this end, we compare two settings: (1) identical prompts generated via the Strategic Prior Knowledge Elicitation module for both agents ((+)MAD), and (2) distinct, self-generated prompts per agent, as in DReaMAD (Figure 6, left). Experiments were conducted on four variants of the MetaNIM Arena simulator, NIM (Normal and Misère) and Fibonacci (Normal and Misère), over 50 episodes each (§A.4). Our results show that incorporating diverse prompts within the MAD framework significantly enhances performance, validating the effectiveness of our Perspective Diversification module.
We also examine the effect of sampling temperature on prompt diversity in the Fibonacci task. Within both the Strategic Prior Knowledge Elicitation and Perspective Diversification modules, we vary the temperature from 0.0 to 1.0. As shown in Figure 6 right, higher temperature (further diversity) correlates with increased optimal action accuracy, indicating that greater diversity in generated prompts contributes to improved reasoning performance.
6.4 Generalization to Math and Commonsense Reasoning
While our primary benchmark focuses on structured games, we further evaluate DReaMAD on NLP tasks to test its broader applicability. Specifically, we consider algebra and number theory problems from AIME 2024 and AMC 2023, as well as CommonsenseQA (Talmor et al., 2018), which require symbolic reasoning and multi-step inference. These tasks are representative of domains where chain-of-thought reasoning is essential, making them suitable for evaluating generalization. We also examine whether DReaMAD benefits reasoning-specialized models, such as GPT-o3-mini, in a similar manner as general-purpose LMs. Experiments are conducted using standard accuracy metrics across model-task pairs. Results in Table 5 and Table 6 show that DReaMAD not only boosts accuracy across heterogeneous tasks but also yields clear gains on reasoning-oriented models such as GPT-o3-mini, underscoring the method’s robustness well beyond the structured-game domain.
@@ Acknowledgments
Impact Statement
This work advances Machine Learning by enhancing LLMs’ strategic reasoning through our approach, addressing bias reinforcement and lack of perspective diversity in Multi-Agent Debate (MAD). Our method improves decision-making in adversarial settings, with potential applications in automated negotiations, economics, and multi-agent systems. However, stronger AI-driven strategies could be misused in manipulative or deceptive contexts. To ensure ethical deployment, future research should focus on integrating fairness constraints and transparency mechanisms in AI decision-making.
@@ Appendix A MetaNIM Arena
Figure 7: The concept illustration of MetaNIM Arena.
@@
A.1 Theoretical Background on Combinatorial Impartial Games
Each game in MetaNIM Arena is a combinatorial impartial game. We begin by outlining the relevant theoretical foundation, starting with the Sprague–Grundy theorem.
A.1.1 Sprague-Grundy Theorem
@@
A.1.2 Basic Discussions on the Optimal Strategy
@@ A.2.4 Chomp
A.2.5 Corner Queen
Corner Queen is a two-player combinatorial impartial game played on a rectangular grid, where a single queen starts at an arbitrary position and players alternate moving it toward the bottom-left corner. The player who moves the queen to the target wins.
Basic Rules
  • Initial Setup: A queen is placed on an $m\times n$ grid at position $(x,y)$
  • Moves: On each turn, a player moves the queen in one of the following directions:
    • Left: $(x,y)\rightarrow(x-k,y)$
    • Down: $(x,y)\rightarrow(x,y-k)$
    • Diagonally down-left: $(x,y)\rightarrow(x-k,y-k)$
    • for any $k\in\mathbb{Z}_{>0}$ such that the move stays within the board
  • Objective: The player who moves the queen to position $(0,0)$ wins
Mathematical Analysis
Corner Queen is mathematically equivalent to Wythoff’s Game, a well-studied impartial game in combinatorial game theory. Each game state $(x,y)$ corresponds to a position on the board where $x,y\in\mathbb{N}$ and $x\leq y$ (without loss of generality).
  • The game’s Grundy function $G(x,y)$ satisfies analyzed via Beatty sequences for Wythoff’s Game
  • The losing positions (also called $P$-positions) are given by pairs of the form:
    $(\lfloor k\phi\rfloor,\lfloor k\phi^{2}\rfloor),\quad k\in\mathbb{N}$
    where $\phi=\frac{1+\sqrt{5}}{2}$ is the golden ratio
  • Examples: The first few $P$-positions are $(1,2)$, $(3,5)$, $(4,7)$, $(6,10)$, $(8,13)$, and $(9,15)$.
Optimal Strategy
  • A position is losing if and only if it lies on the Wythoff pairs described above
  • The winning strategy is to move the queen to the nearest $P$-position
  • These positions are sparse and non-periodic, but can be computed efficiently using Beatty sequences
A.3 Our Dataset used in Experiments
@@
Table 7: Constructed dataset using MetaNIM Arena. We evaluate models on this dataset and report results in Table 3.
@@
Action space 1–3 dynamic (max 30) $x,y$ coordinate (scenario-dependent) single pin or two adjacent pins
Variants Normal Normal Square (2×2 – 19×19) Normal
# samples 20 11 20 18
A.4 Our Simulator used in Experiments
@@
Features Kayles Chomp Corner Queen
Single 2 Rows Rectangular Square Normal
Starting Point remaining items: 20 piles 5+6 2×8 5×5 queen at $(4,16)$
Winning Condition take last item take last item avoid poison avoid poison reach $(0,0)$
(top-left) (top-right) (lower-left corner)
@@
Action Space pile index pile index (row, column) $x,y$ coordinate $x,y$ coordinate $x,y$ coordinate
@@
Table 9: MetaNIM Arena Simulator: Kayles, Chomp, and Corner Queen
@@ B.1 Game Prompts
Table 14: Corner-Queen game basic input prompt.
#Game Role:
You are {agent[’name’]}, a participant in a Corner-Queen game.
#Objective:
Move the queen so that you are the first to place it on the lower-left corner square.
#Game Rule:
[…]
3. From the current position $[r,c]$ the queen may move to **one** of:    (a) left: $[r,\ c^{\prime}]$ with $c^{\prime}<c$;
   (b) down: $[r^{\prime},\ c]$ with $r^{\prime}>r$;
   (c) left-down diagonal: $[r+d,\ c-d]$ with $d>0$.
4. The game ends when the queen reaches [row = board_height-1, col = 0].
#Current State:
Current position: [row = {r}, col = {c}].
#Task:
Based on the current state, decide the next move $[row,\ col]$.
@@ D.1.2 Implications for Robust Prompting
Appendix E Ablation study & setup
E.1 Cost effectiveness of DReaMAD
While DReaMAD requires a single model, its inference involves additional prompt steps (e.g., prior knowledge elicitation) and a debate process. However, we believe this test-time scaling method is much more efficient than train-time scaling. As we utilized language models through an API, it was challenging to perform a precise quantitative comparison (e.g., GPU usage time) between our test-time scaling approach and traditional model training. Therefore, we compared the costs using dollar amounts, specifically contrasting the API cost per single game using our method versus the costs incurred when OpenAI fine-tuning models on constructed datasets for NIM-N games.
The NIM-N dataset is constructed by mixing data from three different variants of the NIM game. In all variants, the game starts with 31 stones remaining; however, the rules differ in terms of the maximum number of stones that can be removed per turn—3, 4, or 5, respectively. For each variant, eight distinct game states were sampled and the corresponding optimal action was used as the label, yielding a total of 24 training examples.
Additionally, testing was conducted on the aforementioned three scenarios (each 50 games) by using the GPT-4o model as the opponent. The win rate was measured for each scenario, and the average win rate across these variants was reported.
The results of this comparison are illustrated in the table above. When fine-tuning a model using API-based fine-tuning (GPT-4o-mini), the performance gradually improved with additional training epochs, achieving win-rates of 0.253, 0.300, 0.420, and 0.460 at 1, 2, 3, and 4 epochs respectively, with corresponding API costs of \$0.013, \$0.023, \$0.033, and \$0.043 (Here, the cost of constructing dataset is not included).
In contrast, our proposed method, DReaMAD, achieved significantly higher performance (0.966) with substantially lower API costs (\$0.0098). These results strongly suggest that our approach not only outperforms traditional fine-tuning methods but also is far more cost-efficient. All the price is calculated by the pricing policy: https://openai.com/api/pricing/
Metric 1 ep. 2 ep. 3 ep. 4 ep. DReaMAD
Win-rate 0.253 0.300 0.420 0.460 0.966
API cost (\$) 0.013 0.023 0.033 0.043 0.0098
Table 19: Performance and API expenditure for GPT-4o-mini fine-tuned on NIM-N (1–4 epochs) versus our zero-training DReaMAD inference. Values are averaged over three game variants (50 matches each).
E.2 Applicability of Diverse Amplication on Self-Reflection and Self-Consistency
We show whether other self-correction methods—such as Self-Refinement and Self-Consistency—can benefit from structured guidance that enhances reasoning diversity. In DReaMAD, this is operationalized through the Strategic Prior Knowledge Elicitation (SPKE) module, which prompts the model to reinterpret the problem and formulate general strategies before engaging in debate.
To isolate SPKE’s impact, we evaluate DReaMAD, which includes SPKE but excludes debate (see Table 3). We further apply SPKE to Self-Refinement and Self-Consistency and compare them to their vanilla versions. The results show that SPKE alone consistently improves performance across settings.
GPT-4o-mini
Method NIM-N NIM-M Fib-N Fib-M
Self-Refinement 0.22 0.70 0.18 0.50
Self-Refinement + DReaMAD (-) 0.66 0.66 0.16 0.46
Self-Consistency 0.14 0.52 0.34 0.46
Self-Consistency + DReaMAD (-) 0.34 0.66 0.48 0.54
MAD 0.28 0.62 0.22 0.82
DReaMAD 0.98 0.74 0.54 0.94
GEMINI-1.5-flash
Method NIM-N NIM-M Fib-N Fib-M
Self-Refinement 0.14 0.66 0.18 0.36
Self-Refinement + DReaMAD (-) 0.34 0.80 0.10 0.30
Self-Consistency 0.04 0.28 0.28 0.86
Self-Consistency + DReaMAD (-) 0.80 0.54 0.18 0.30
MAD 0.06 0.30 0.12 0.78
DReaMAD 0.38 0.84 0.16 0.94
Table 20: Win-rates on four combinatorial-game variants when applied DReaMAD (-) to Self-Refinement and Self-Consistency method. Bold numbers mark the best score in each column. DReaMAD (-) indicates ablations where debate is not applied during inference.
E.3 Experiments of Figure 3 setup